Please complete all parts
My GradesBookmarksGmailUTSA Self-Development YouTube Feedly SR Qual Company builds snowboards. Qualil Company has reported the following costs for the previous year. Assume no production inventories en r tergse nera 85,000 S 42,000 $130,000 5 20,000 S 81,000 S 45,000 34000 17,000 27.000 23,000 $ 1.300 80,000 S 46.000 Factory rent Uliies for factory Wages of assembly workers Production supervisor salary Sales manager salary Factory property taxes Depreciation on production equipment Cost of bindings (newrial Screws Wages of snowboand painters Wages for maintenance workers a. Compute the direct material costs. ?, computere dret labor cost. Computn the marufacturing overhead esc 7 lab
media%2Faf0%2Faf01b3a2-7946-4f89-a56a-23
Answer #1

Answer:

a. Direct material costs

108000

b. Direct labor cost

161000

c. Manufacturing overhead

576300

d. Total manufacturing cost

845300

e. Prime cost

269000

f. Conversion cost

737300

g. Total period cost

164000

Working notes for the above answer is as under

a. Direct material costs

=85000+23000

=108,000

___________________________
b. Direct labor cost

=81000+80,000

=161,000

_______________________________
c. Manufacturing overhead

=420000+20,000+45000+17000+27000+1300+46000

=$576300

__________________________________

d. Total manufacturing cost
=108,000+161,000+576,300

=845,300

___________________________________

e. Prime cost

=108000+161000

=269,000

________________________________
f. Conversion cost

=161000+576300

=$737,300

___________________________________
g. Total period cost

=130,000+34,000

=164,000